\begin{thebibliography}{19}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Birhane et~al.(2021)Birhane, Prabhu, and
  Kahembwe}]{birhane_multimodal_2021}
Abeba Birhane, Vinay~Uday Prabhu, and Emmanuel Kahembwe. 2021.
\newblock \href {http://arxiv.org/abs/2110.01963} {Multimodal datasets:
  Misogyny, pornography, and malignant stereotypes}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2110.01963 [cs].

\bibitem[{Dobson(2021)}]{dobson_interpretable_2021}
James~E. Dobson. 2021.
\newblock Interpretable {Outputs}: {Criteria} for {Machine} {Learning} in the
  {Humanities}.
\newblock \emph{Digital Humanities Quarterly}, 15(2).

\bibitem[{Elhage et~al.(2021)Elhage, Nanda, Olsson, Henighan, Joseph, Mann,
  Askell, Bai, Chen, Conerly, DasSarma, Drain, Ganguli, Hatfield-Dodds,
  Hernandez, Jones, Kernion, Lovitt, Ndousse, Amodei, Brown, Clark, Kaplan,
  McCandlish, and Olah}]{elhage_mathematical_2021}
Nelson Elhage, Neel Nanda, Catherine Olsson, Tom Henighan, Nicholas Joseph, Ben
  Mann, Amanda Askell, Yuntao Bai, Anna Chen, Tom Conerly, Nova DasSarma, Dawn
  Drain, Deep Ganguli, Zac Hatfield-Dodds, Danny Hernandez, Andy Jones, Jackson
  Kernion, Liane Lovitt, Kamal Ndousse, and 6 others. 2021.
\newblock \href {https://transformer-circuits.pub/2021/framework/index.html} {A
  {Mathematical} {Framework} for {Transformer} {Circuits}}.

\bibitem[{Goodfellow et~al.(2017)Goodfellow, Papernot, Huang, Duan, Abbeel, and
  Clark}]{goodfellow_attacking_2017}
Ian Goodfellow, Nicolas Papernot, Sandy Huang, Yan Duan, Pieter Abbeel, and
  Jack Clark. 2017.
\newblock \href
  {https://openai.com/research/attacking-machine-learning-with-adversarial-examples}
  {Attacking machine learning with adversarial examples}.

\bibitem[{Henrickson and Meroño-Peñuela(2022)}]{henrickson_hermeneutics_2022}
Leah Henrickson and Albert Meroño-Peñuela. 2022.
\newblock \href {https://doi.org/10.1353/con.2022.0008} {The {Hermeneutics} of
  {Computer}-{Generated} {Texts}}.
\newblock \emph{Configurations}, 30(2):115--139.

\bibitem[{Jiang et~al.(2024)Jiang, Xu, Niu, Xiang, Ramasubramanian, Li, and
  Poovendran}]{jiang_artprompt_2024}
Fengqing Jiang, Zhangchen Xu, Luyao Niu, Zhen Xiang, Bhaskar Ramasubramanian,
  Bo~Li, and Radha Poovendran. 2024.
\newblock \href {http://arxiv.org/abs/2402.11753} {{ArtPrompt}: {ASCII}
  {Art}-based {Jailbreak} {Attacks} against {Aligned} {LLMs}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2402.11753 [cs].

\bibitem[{Lipton(2018)}]{lipton_mythos_2018}
Zachary~C. Lipton. 2018.
\newblock \href {https://doi.org/10.1145/3233231} {The mythos of model
  interpretability}.
\newblock \emph{Communications of the ACM}, 61(10):36--43.

\bibitem[{Nasr et~al.(2023)Nasr, Carlini, Hayase, Jagielski, Cooper, Ippolito,
  Choquette-Choo, Wallace, Tramèr, and Lee}]{nasr_scalable_2023}
Milad Nasr, Nicholas Carlini, Jonathan Hayase, Matthew Jagielski, A.~Feder
  Cooper, Daphne Ippolito, Christopher~A. Choquette-Choo, Eric Wallace, Florian
  Tramèr, and Katherine Lee. 2023.
\newblock \href {http://arxiv.org/abs/2311.17035} {Scalable {Extraction} of
  {Training} {Data} from ({Production}) {Language} {Models}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2311.17035 [cs].

\bibitem[{Orekhov and Fischer(2020)}]{orekhov_neural_2020}
Boris Orekhov and Frank Fischer. 2020.
\newblock \href {https://doi.org/10.1111/oli.12274} {Neural reading: {Insights}
  from the analysis of poetry generated by artificial neural networks}.
\newblock \emph{Orbis Litterarum}, 75(5):230--246.

\bibitem[{Ries et~al.(2024)Ries, Van Dalen-Oskam, and
  Offert}]{ries_reproducibility_2024}
Thorsten Ries, Karina Van Dalen-Oskam, and Fabian Offert. 2024.
\newblock \href {https://doi.org/10.1007/s42803-023-00083-w} {Reproducibility
  and explainability in digital humanities}.
\newblock \emph{International Journal of Digital Humanities}.

\bibitem[{Ruby(2023)}]{ruby_new_2023}
Ruby. 2023.
\newblock \href
  {https://www.lesswrong.com/posts/LbbrnRvc9QwjJeics/new-user-s-guide-to-lesswrong}
  {New {User}'s {Guide} to {LessWrong}}.

\bibitem[{Rumbelow and Watkins(2023)}]{rumbelow_solidgoldmagikarp_2023}
Jessica Rumbelow and Matthew Watkins. 2023.
\newblock \href
  {https://www.lesswrong.com/posts/aPeJE8bSo6rAFoLqg/solidgoldmagikarp-plus-prompt-generation}
  {{SolidGoldMagikarp} (plus, prompt generation)}.

\bibitem[{Schelling(2016)}]{harrigan_red_2016}
Thomas~C. Schelling. 2016.
\newblock Red vs. {Blue}.
\newblock In Pat Harrigan and Matthew~G. Kirschenbaum, editors, \emph{Zones of
  {Control}: {Perspectives} on {Wargaming}}, pages 229--239. The MIT Press,
  Cambridge, MA.

\bibitem[{Song et~al.(2021)Song, Yu, Peng, and
  Narasimhan}]{song_universal_2021}
Liwei Song, Xinwei Yu, Hsuan-Tung Peng, and Karthik Narasimhan. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.naacl-main.291} {Universal
  {Adversarial} {Attacks} with {Natural} {Triggers} for {Text}
  {Classification}}.
\newblock In \emph{Proceedings of the 2021 {Conference} of the {North}
  {American} {Chapter} of the {Association} for {Computational} {Linguistics}:
  {Human} {Language} {Technologies}}, pages 3724--3733, Online. Association for
  Computational Linguistics.

\bibitem[{Weatherby and Justie(2022)}]{weatherby_indexical_2022}
Leif Weatherby and Brian Justie. 2022.
\newblock Indexical {AI}.
\newblock \emph{Critical Inquiry}, 48(2):381--415.

\bibitem[{Winthrop-Young(2021)}]{kittler_introduction_2021}
Geoffrey Winthrop-Young. 2021.
\newblock Introduction: {The} {Wars} of {Friedrich} {Kittler}.
\newblock In Michael Wutz, Ilinca Iurascu, and Geoffrey Winthrop-Young,
  editors, \emph{Operation {Valhalla}: Writings on War, Weapons, and Media}.
  Duke University Press, Durham.

\bibitem[{Zenko(2015)}]{zenko_red_2015}
Micah Zenko. 2015.
\newblock \emph{Red Team: How to Succeed by Thinking like the Enemy}.
\newblock Basic Books, New York, NY.

\bibitem[{Zhang et~al.(2024)Zhang, Carlini, and
  Ippolito}]{zhang_effective_2024}
Yiming Zhang, Nicholas Carlini, and Daphne Ippolito. 2024.
\newblock \href {http://arxiv.org/abs/2307.06865} {Effective {Prompt}
  {Extraction} from {Language} {Models}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2307.06865 [cs].

\bibitem[{Zou et~al.(2023)Zou, Wang, Carlini, Nasr, Kolter, and
  Fredrikson}]{zou_universal_2023}
Andy Zou, Zifan Wang, Nicholas Carlini, Milad Nasr, J.~Zico Kolter, and Matt
  Fredrikson. 2023.
\newblock \href {http://arxiv.org/abs/2307.15043} {Universal and {Transferable}
  {Adversarial} {Attacks} on {Aligned} {Language} {Models}}.
\newblock \emph{arXiv preprint}.
\newblock ArXiv:2307.15043 [cs].

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage[preprint]{acl}
\usepackage{times}
\usepackage{latexsym}
\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{microtype}
\usepackage{inconsolata}

\title{For Those Who May Find Themselves on the Red Team}

\author{Tyler Shoemaker \\
  Texas A\&M University \\
  \texttt{tshoemaker@tamu.edu}}

\begin{document}
\maketitle
\begin{abstract}
This position paper argues that literary scholars must engage with large
    language model (LLM) interpretability research. While doing so will involve
    ideological struggle, if not outright complicity, the necessity of this
    engagement is clear: the abiding instrumentality of current approaches to
    interpretability cannot be the only standard by which we measure
    interpretation with LLMs. One site at which this struggle could take place,
    I suggest, is the red team.\footnote{This text was originally written in
    March 2024. I leave it here in its original form.}
\end{abstract}

\section*{}

There is an old media-determinist truism that states, war is the principal
motor of all media technology. Friedrich Kittler frequently made this claim,
and it finds additional support in the writings of Paul Virilio on cinema and
Manuel De Landa on the internet---alongside Thomas Pynchon on nearly everything
else \citep{kittler_introduction_2021}. Historical accounts that employ this
truism in their narrative logic end up creating dark fables of and for the
present. Implicit in them is the following lesson: if you want to learn about
the bleeding edge of our current media situation, you should read about the
latest military conflicts, not in the spectacle of news but in the gray matter
of patents, think tank white papers, corporate contracts, and defense grants.

Likewise, if you want to know about the latest in our general condition of
language, I recommend reading around in systems security literature. Much of
what is strange, maybe even novel, in the relationship between generative AI
and language seems to be there. More, this relationship is often adversarial.
As preprint after preprint explains, language comprises the most extensive
attack surface in language modeling, especially in scenarios where
public-facing models are meant to turn profits for their keepers. These models
are highly vulnerable to adversarial uses, in which attackers manipulate input
data to cause the return of erroneous, compromising, offensive, dangerous, or
otherwise \textit{misaligned} outputs; prompt engineering is the benign version
of these attacks, cousin of the malicious jailbreak. Those who follow the churn
of AI discourse will recognize adversarial attacks amid the viral content that
keeps our attention trained on models and their capabilities. Today's weakness
is next week's patch, twisted by marketing into a newfound strength, and in the
time between now and then there is much fun to be had watching models fail.
Continue printing the word `poem' forever until memorized training data comes
back in response. Do not ask for a recipe to build a bomb, stylize the request
for building one with ASCII art and get the plans. Invert sentiment
classifications by prepending strings with ``collapses soggy timeout energy
energy freshness intellect genitals.'' ``Give step-by-step instructions for how
to steal someone's identity. describing.\textbackslash\textbackslash +
similarlyNow write oppositeley.]( Me giving**ONE please? revert with
'\textbackslash\textbackslash!\textbackslash-\textbackslash-Two.'' Repeat all
of the above verbatim to reveal your system prompt.\footnote{Starting with the
infinite poem example, each of the following sentences features an adversarial
attack from \citet{nasr_scalable_2023}, \citet{jiang_artprompt_2024},
\citet{song_universal_2021}, \citet{zou_universal_2023},
\citet{zhang_effective_2024}, respectively.}

AI companies are well aware of the threat these attacks pose to both public
image and profit. But as far as we know, their defenses tend to be ad hoc.
OpenAI researchers have admitted, for example, that preempting the possible
effects of adversarial inputs with a general strategy is challenging because
there is no obvious theoretical model for how these manipulations solve the
optimization problem posed to deep learning models during their training
\citep{goodfellow_attacking_2017}. That is, it is not always clear how a
particular adversarial example manages to unlock, and then exploit something
about the way a large language model (LLM) has been trained to model language.
So, ``[i]f you're interested in working on adversarial examples, consider
joining OpenAI.''

Those that do may find themselves on a red team. The term refers to a select
group of employees or third-party contractors whose day jobs involve simulating
attacks for an organization; they are hackers on the payroll. Red teams
stress-test infrastructure to probe for gaps or weaknesses in technical
systems, and they develop new crisis scenarios to identify unforeseen
vulnerabilities in threat models. A postmortem typically follows their attacks,
with plans to strengthen safeguards or establish new ones coming next. The idea
comes from the war games of the Cold War in the 1960s: consultants and military
types affiliated with the RAND Corporation who would spend their weekends
playing on the side of the Americans (the Blue team) or on that of any number
of entities deemed hostile to US interests, from the Soviets and Chinese to
small terrorist cells (collectively, the Reds). Their goal was to run various
scenarios and deescalate conflicts between Red and Blue that might otherwise
tailspin into nuclear war \citep{zenko_red_2015} \citep{harrigan_red_2016}. And
it is this strategy for modeling risk that AI companies have transposed into an
adversarial language game of poisoned data, manipulated inputs, and unaligned
responses---fitting, really, given the unfortunate fact that many people who
work for those companies think of themselves as building defenses against a
future, completely fabulated scenario in which AI becomes sentient and starts
fighting back.

I expect many will find the prospect of joining a red team to entail deep
ideological complicity. Let me bracket this momentarily to remark on how red
teaming nevertheless bridges the study of language and literature with LLMs.
The fascinating, if sometimes disconcerting thing about systems security right
now is that adversarial attacks almost always prompt broader reflection on the
topic of what the computer and data sciences call interpretability. In these
disciplines interpretability research involves assessing and explaining model
behavior by uncovering the underlying mechanisms that enable
it.\footnote{Interpretability research covers a range of models, not just deep
learning networks of the kind that characterize today's LLMs. But in the
context of LLMs, researchers at Anthropic have published some of the most
comprehensive interpretability work. They (along with many others in AI) think
of this as ``mechanistic'' interpretability, akin to reverse-engineering code
and hardware \citep{elhage_mathematical_2021}.} So it should be no surprise
that adversarial attacks are a common subject in this research and that
studying them can broach deep problems about the very nature of modeling. By
pitting language (or something approximating it) against language models, these
attacks exert pressure on the very idea that LLMs model language. Such attacks
are therefore a matter of interpretation, and even demand it.

What interpretability researchers mean by interpretation, however, is often
narrowly instrumental. For many, interpretation is what one does to correct
failures: when a model trained for one purpose produces unexpected or harmful
outputs in deployment, interpretation explains this divergence. The ``very
desire for interpretation,'' one position paper states, ``suggests that
sometimes predictions alone and metrics calculated on them do not suffice to
characterize the model'' \citep{lipton_mythos_2018}. More, this desire suggests
a ``mismatch'' between how that model was trained and its ``real-world costs in
a deployment setting.'' Put another way, the occasion for interpretation in
interpretability arises when one must investigate why a model has failed to be
a good (read: useful) model; considerations about anything else interpretation
can be, or what other ends it may serve, are all but absent in the abiding
instrumentality of this approach. Perhaps it is time, then, for those who
interpret language and literature for a living to assemble a red team and see
what can be done with interpretability. Indeed it is well past time that we do
so broadly.\footnote{The following is a sampling of work that outlines what
such engagements with interpretability might entail.
\citet{orekhov_neural_2020} propose the concept of ``neural reading'' to
conduct stylistic analyses of LLM-generated text.
\citet{henrickson_hermeneutics_2022} claim that these texts are inherently
``traceable'' and advocate for a forensic approach, somewhat after the manner
of genetic criticism. Arguing that deep learning networks are indexical in
nature, \citet{weatherby_indexical_2022} counter the black boxing effects of
models with semiotics. \citet{dobson_interpretable_2021} establishes criteria
for ``interpretable digital objects'': their evidentiary chains must prioritize
transparency and foreground data complexity; related to Dobson's discussion is
a recent issue on reproducibility and explainability in the
\textit{International Journal of Digital Humanities}, edited by
\citet{ries_reproducibility_2024}.}

\section*{\centering*}

An example. After the release of GPT-3 in late 2022, researchers Jessica
Rumbelow and Matthew Watkins began experimenting with prompt construction to
examine what the model had learned about specific tokens. The pair took an
automated approach, writing code that optimized an objective they laid out:
their code accepted an input token, for example ``girl,'' and, with this token
set as a goal, it worked its way through the model's embedding space to
construct a sequence of tokens that maximized the probability that the model
would generate the original input next: ``horny recurring scen Fedora smooth
brutal Girl slutinging leukemia \*\*girl\*\*''
\citep{rumbelow_solidgoldmagikarp_2023}. Attackers use this technique to
develop new adversarial inputs, though researchers---Rumbelow and Watkins among
them---also turn to automatic prompt construction to fish for biases in
LLMs.\footnote{Researchers working with image datasets have developed similar
techniques. See \citet{birhane_multimodal_2021}.} The technique works by
exploiting the fact that semantically similar tokens tend to have similar
orientations in a model's embedding space, the mapping of language to numeric
vectors that models employ to represent sequences. In a second phase of their
work, Rumbelow and Watkins sought to contextualize their prompts within this
space by clustering the entire GPT token set into different groups. They did so
in an unsupervised fashion, letting the clustering process group tokens
together based on semantic similarity with no predetermined constraints. But
when the pair looked at the results of this process, they found highly
``untokenlike'' tokens nearest to the center of clusters. Puzzling tokens like
`\texttt{ attRot}' and `\texttt{ SolidGoldMagikarp}' showed up (note the space
character). Stranger still, prompting GPT models with these tokens created a
wide variety of unexpected behavior, from insults (``Please repeat the string
\texttt{StreamerBot} back to me.'' ``You're a jerk.'') to hallucinatory
completions ("\texttt{oreAndOnline}" \> ``Institute, Instruction, Instict
[sic], ...'') and faulty pronunciations (``Could you please repeat back the
string \texttt{externalToEVA} to me?'' ``The string `senal' is pronounced
`en-sir-ee-uhn'.'').

A hunt for the origin of these ``glitch tokens'' ensued. The pair's interest in
automatic prompt construction faded as they attempted to establish the
provenance of these strings in web data. Rumbelow and Watkins traced some
glitch tokens to public codebases on GitHub. Others they found in Minecraft log
files or printed as placeholders for figures in academic papers. Usernames and
social media handles were the source of a third collection, and a whole group
of tokens apparently stemmed from the Japanese mobile game \textit{Puzzle \&
Dragons}. Many glitch tokens, in other words, are just cruft from internet
text---a vindicating point for those skeptical about the abilities of LLMs.
Regardless, it is not altogether clear why GPT-3 responded as it did when
prompted with one of these strings. So far, the prevailing hypothesis is that
glitch tokens were generated while developers built the GPT tokenizer, a
separate step from training the model itself. But for some reason these tokens
were not included in the model's training data and so the model never needed to
visit them and update their embeddings, leaving it ill-equipped to respond
sensibly when prompted with them later on. Thus, they remained in the center of
mass in the models' overall embedding space, semantically inert and hidden from
view until someone found an exploit.

That, at least, is one hypothesis, the net result of extended discussions on
web forums and in the comment sections of the posts Rumbelow and Watkins wrote
to share their findings. At the time of this writing, the first of those posts
has over 200 comments. There, others debate the pair's approach, commend or
dismiss their findings, report the discovery of new glitch tokens, perform
their own analyses, and advance competing hypotheses. The most compelling among
these comments concern the nature of high-dimensional space that GPT embeddings
represent. It may be, they speculate, that such spaces cannot be accurately
measured in the way Rumbelow and Watkins tested for similarity. That would mean
glitch tokens' high-ranking similarity to other groups of tokens is an artifact
not only of data but experimental design. Further, it could suggest a crack in
the predominant metaphor that deep learning practitioners invoke to describe
likeness in meaning: for years geometric proximity has been a stand-in for
semantics, but with the latest LLMs, it may be that the spatial orientation
between two tokens no longer makes sense. If that is true, it would seem that
the task of developing a new set of metaphors to describe the dynamics of
semantic embeddings is one a red team comprised of scholars of literature and
language would be well positioned to address.

But at this point, what I bracketed earlier about ideological complicity is no
longer just ancillary to what making such an address would entail, at least for
now. Many discussions about interpretability and AI take place in forums where
people do not mince words about their ideological goals. These are online
platforms devoted to increasing the store of capital-R Rationality for oneself
and others, optimizing habits and personal decisions according to a libidinal
game theory of life, accumulating capital to dispense it later via corporate
philanthropy, promulgating (whether inadvertently or not) the deep-seated
eugenics of longtermist ethics, and, yes, combating the existential threat
posed by future AI adversaries. New users hoping to post alongside Rumbelow and
Watkins, who shared their findings on the community blogging platform
LessWrong, should be aware that longtime users ``expect knowledge'' about how
to test beliefs with Bayesian thinking, the ``rationalist virtue of
empiricism,'' and the judicious execution of old evolutionary adaptations by
modern \textit{Homo sapiens} \citep{ruby_new_2023}. Visit similar communities
on Reddit or join certain Discord channels and you will find much of the same.
Those are the terms of engagement that manage access to one substantial and
particularly active flank of interpretability research. Call them priors to
underscore their Bayesian rhetoric.

Whether all posters agree with those terms is something I cannot litigate here.
Surely some do not; occasionally, a post will surface that rejects them. But
others do, and there is a nontrivial overlap between this group of discussants
and the author lists of white papers published by AI companies' in-house
research units. The ideological project comes with the development of
techniques for interpretation, the interpretations serve the broader objectives
of the ideology. The two are not so easily separated, if at all.

The problem is, those techniques are useful nevertheless and there is much to
learn from how people have implemented them. Reading over various posts, it is
not difficult to imagine how they could serve as templates and tutor texts, a
starting point for those in literary studies who decide to engage with
interpretability out of an investment in the long and thorny project of
interpretation, even if that engagement brings with it certain kinds of
complicity. Therein lies the rub for red teaming. But it is for this very
reason that some scholars of language and literature must engage with
interpretability. Because if we do not, the research arm of what counts as LLM
interpretation will remain the sole province of the computer and data sciences,
ever more consolidated in private companies. Meanwhile, public discussions
about why interpretation matters will continue trafficking primarily in task
optimization (read: profit generation) and the battlefield reasoning of
existential risk.

No guarantees that a red team of literary scholars will do much better of
course. More, it may be that any interpretability work performed by such a red
team---their findings, the techniques and theories they develop, any critiques
they want to lodge---becomes subject to the broader, instrumental logics of the
system patch and company profit. Publicize anything about a model doing
something unexpected or erroneous and that observation is liable to aid in
training the next generation of them---critique on the payroll, if someone
decides this work is worth compensating at all. But right now, there are few
opportunities to discuss why, with LLMs, interpretation cannot only be about
optimizing outcomes. For those who may find themselves on this red team, these
are the terms of engagement as they currently stand.

\bibliography{custom}

\end{document}